OpenStudy (anonymous):

Find the horizontal intercepts of the quadratic by rewriting the quadratic into transformation form. f(x) = -4x^2+10x-5

hero (hero):

@jtryon, are you here?

OpenStudy (anonymous):

yes, I am here

hero (hero):

You are given f(x) = -4x^2 + 10x - 5 and you want to write the quadratic in transformational form, correct?

OpenStudy (anonymous):

correct, now I know that one way to find the x-intercepts is to use the quadratic formula but I saw in my textbook that you can also find the x-int. once you write the quadratic in transformational form but I am not understanding that part

hero (hero):

The standard form of a quadratic equation is \(y = a(x - h)^2 + k\) and the transformational form is \(\dfrac{1}{a}(y - k) = (x - h)^2\) correct?

OpenStudy (anonymous):

correct

hero (hero):

Which means that if we convert the given quadratic to standard form, it should be relatively easy to convert it to transformational form, right?

OpenStudy (anonymous):

no, wait

OpenStudy (anonymous):

the standard form is defined as ax^2+bx+c and the transformational form is a(x-h)^2+k

OpenStudy (anonymous):

yes, converting it to transformational form was relatively easy, but with the vertex that I found at 5/4, it was then difficult to plug that back in and come up with the x-intercepts

OpenStudy (anonymous):

The example that they have given use a whole integer vertex of -1 so that is easy to plug in

hero (hero):

What did you get for the transformational form of f(x) = -4x^2+10x-5 ?

OpenStudy (anonymous):

f(x) = -4(x+1)^2+1.25

OpenStudy (anonymous):

I think that is where I may have made and error

OpenStudy (anonymous):

@Hero

hero (hero):

Would you mind showing your steps please?

OpenStudy (anonymous):

Sure -4(5/4)^2+10(5/4) - 5 -6.25 + 12.5 - 5 = 1.25

OpenStudy (anonymous):

I am not certain about the (x+1) in the transformational form as I still have only a slight understanding of this form

OpenStudy (anonymous):

If I got the idea of transformational form, I''m thinking it would be easier to find the x-intercepts than having to use the quadratic formula

hero (hero):

I got something different for the transformational form

hero (hero):

@jtryon

OpenStudy (anonymous):

What did you get? I probably made a mistake somewhere

OpenStudy (anonymous):

The first step I did was solve for my h which is 1.25 and then my k which is also 1.25

hero (hero):

There are a specific set of steps that you perform to convert from standard form to vertex form.

hero (hero):

The transformational form that they are describing here is also known as the vertex form of a quadratic equation.

hero (hero):

First, @jtryon, since y = f(x), we can write the given quadratic equation as \(y = -4x^2 + 10x - 5\)

OpenStudy (anonymous):

Yes, that is what I like to do instead of have it as f(x) or h(r)

OpenStudy (anonymous):

It helps me to call it y

hero (hero):

Next what we can do is multiply (or divide) both sides by -1 to get \(-y = 4x^2 - 10x + 5\) Do you agree?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Why do we want to change it to -y? I understand the process of multiplying both sides by -1

hero (hero):

You'll see... Afterwards, we can factor 4 from the quadratic expression to get \(-y = 4(x^2 - \dfrac{10}{4}x + \dfrac{5}{4})\) agreed?

OpenStudy (anonymous):

Ok, continue

hero (hero):

Now we divide both sides by 4 to get \(-\dfrac{y}{4} = x^2 - \dfrac{10}{4}x + \dfrac{5}{4}\)

hero (hero):

Since \(\dfrac{10}{4} = \dfrac{5}{2}\) we can re-write the previous step as: \(-\dfrac{y}{4} = x^2 -\dfrac{5}{2}x + \dfrac{5}{4}\) Agreed?

OpenStudy (anonymous):

That makes sense

hero (hero):

Now we will subtract \(\dfrac{5}{4}\) from both sides to get \(-\dfrac{y}{4} - \dfrac{5}{4} = x^2 - \dfrac{5}{2}x\) Which leads us to the whole point of doing all that manipulation. Notice that we have the form \(x^2 + bx\) on the right side. Do you know what this means?

OpenStudy (anonymous):

That is the standard form?

hero (hero):

Not quite. But if we add \(\left(\frac{b}{2}\right)^2\) to both sides, we'll end up with a complete square on the right side. Does the concept of "completing the square" sound familiar to you?

OpenStudy (anonymous):

Yes, it is familiar to me. I have not used that in awhile.

OpenStudy (anonymous):

How do we get the x-intercepts when we have changed from standard form to transformational form?

hero (hero):

Once we finish converting to transformational/vertex form, we'll be able to find these x-intercepts you refer to.

hero (hero):

The interesting thing is that converting to tranformational form is the complete long way to find x-intercepts.

OpenStudy (anonymous):

There is a series of steps you have to complete before you convert to transformational form so I am wondering if it is faster to use the quadratic formula

OpenStudy (anonymous):

It is relatively easy to identify a,b,c from the terms in the quadratic

OpenStudy (anonymous):

Oh, so it is a long way of finding the x-intercepts

hero (hero):

Basically, if your goal was to find the x-intercepts, converting from general form to transformational form is about the most time consuming possible method to use. Especially considering this quadratic.

OpenStudy (anonymous):

They made it seem really simple in the textbook using the following example: y = 2x^2+4x-4 0 = 2x^2 + 4x - 4 h = b/2a = 4/2(2) = -1 k = f(-1) = 2(-1)^2+4(-1)-4=6 0 = 2(x+1)^2-6 6 = 2(x+1)^2 3 = (x+1)^2 x+1=+- sqrt(3) x = -1+- sqrt(3)

OpenStudy (anonymous):

It wasn't that simple when I attempted to use it with the vertex of (1.25, 1.25)

hero (hero):

Scan the page from your textbook and upload it here so I can see exactly how they typed out the steps.

OpenStudy (anonymous):

@Hero

hero (hero):

But yeah, they are right. The method to find (h,k) is much faster than completing the square.

hero (hero):

Because (h,k) is the vertex of the parabola. So actually, it is silly to complete the square.

hero (hero):

Basically the vertex (transformational) form of a quadratic is \(y = (x - h)^2 + k\) \(h = -\dfrac{b}{2a}\) \(k = c - \dfrac{b^2}{4a}\) You are given \(f(x) = -4x^2 + 10x - 5\) And a = -4, b = 10, c = -5 First solve for h: \(h = -\dfrac{10}{2(-4)} = \dfrac{10}{8} = \dfrac{5}{4}\) Then solve for k: \(k = -5 - \dfrac{10^2}{4(-4)} = -5 + \dfrac{100}{16} = -5 + \dfrac{25}{4} = -\dfrac{20}{4} + \dfrac{25}{4} = \dfrac{5}{4}\) Therefore the transformational form of f(x) = -4x^2 + 10x - 5 is \(f(x) = -4\left(x - \dfrac{5}{4}\right)^2 + \dfrac{5}{4}\) Check: http://www.wolframalpha.com/input/?i=-4x%5E2+%2B+10x+-+5+%3D+-4%28x+-+5%2F4%29%5E2+%2B+5%2F4